HIGH RESOLUTION MID-INFRARED SPECTROSCOPY OF D_{2}O^{+}

\underline{CHUANXI DUAN}, RUI ZHENG, SONG LI, and GUANGMING HUANG, College of Physical Science and Technology, HuaZhong Normal University, Wuhan 430079, China.

We have observed the mid-infrared absorption spectra of D_{2}O^{+} using a tunable lead-salt diode laser spectrometer with velocity-modulation technique. The D_{2}O^{+} ions were generated in a water-cooled discharge of a mixture of D_{2}O and helium. Spectra were collected in the region between 1000 and 1150 cm^{−1}. Totally 64 lines were assigned to the ν_{2} fundamental band of D_{2}O^{+} (X̃^{2}B_{1}). Precise molecular constants have been determined.